OpenStudy (erniellyd28):

the combined land area of the countries a and b is 132,973 square kilometers. country a is larger by 573 square kilometers. determine the land area of each country

jimthompson5910 (jim_thompson5910):

`the combined land area of the countries a and b is 132,973 square kilometers` so, a+b = 132973

jimthompson5910 (jim_thompson5910):

`country a is larger by 573 square kilometers` meaning that (area of country A) = (area of country B) + 573 a = b + 573

OpenStudy (mathmale):

Please choose variables, such as a and b, or A and B, to represent the respective areas of these two countries. Then write the relevant equations relating these variables to each other.

jimthompson5910 (jim_thompson5910):

the system of equations is therefore a+b = 132973 a = b + 573

OpenStudy (erniellyd28):

thats the answer jim thompson?

OpenStudy (skullpatrol):

You need to solve for a and b :-)

OpenStudy (erniellyd28):

thats where i get lost lol

OpenStudy (erniellyd28):

im very sorry

OpenStudy (skullpatrol):

Replace a in the first equation: a + b = 132, 973 with b + 573, because a = b + 573

OpenStudy (skullpatrol):

a + b = 132, 973 b + 573 + b = 132, 973

OpenStudy (erniellyd28):

do i divide next at all?

OpenStudy (skullpatrol):

Now subtract 573 from both sides, simplify, then divide both sides by 2. This will give you the value for b.

OpenStudy (skullpatrol):

Got it? Then substitute back to get the value for a.

OpenStudy (erniellyd28):

i got 66200

OpenStudy (erniellyd28):

but i dont know if i did it right

OpenStudy (erniellyd28):

or if i missed a step